OpenStudy (hanner_b_nanner):

Several questions, please help...Fan & Medal :) Suppose that the functions f and g are defined for all real numbers as follows. f(x)=4x^2 g(x)=3x Write the expressions for (f*g)(x) and (f+g)(x) and evaluate (f-g)(-2).

RhondaSommer (rhondasommer):

what do you think

OpenStudy (kashmoneyjr):

To evaluate (f*g)(x), we simply multiply the functions: (f*g)(x)=f(x)*g(x) 4x^2*3x=12x^3

jimthompson5910 (jim_thompson5910):

(f*g)(x) is equivalent to f(x)*g(x). These are two ways to say the same thing: multiply the functions f and g. Here's one example Let f(x) = 2x^3 g(x) = 9x (f*g)(x) = f(x)*g(x) (f*g)(x) = 2x^3*9x (f*g)(x) = (2*9)*(x^3*x) (f*g)(x) = (2*9)*(x^3*x^1) (f*g)(x) = 18*x^(3+1) (f*g)(x) = 18*x^4 This is an example and not one of your answers. Hopefully it will help you get the answer you need.

OpenStudy (kashmoneyjr):

the same principle applies to (f+g)(x), and (f-g)(2), except for the last one you substitute in -2

OpenStudy (hanner_b_nanner):

Thank y'all :)